OpenStudy (acacia):

Which of the following is not a geometric sequence? A. 1, 2, 4, 8 B. 3, 1, -1,-3 C. ½, ¼, 1/8, 1/16 D. 216, 72, 24, 8

OpenStudy (mathstudent55):

Do you know what to do to find out?

OpenStudy (acacia):

@jim_thompson5910 can you help me?

OpenStudy (mathstudent55):

Divide the first number by the second, the second by the third, the third by the fourth. If all divisions are equal, it's geometric.

OpenStudy (acacia):

oh okay

OpenStudy (mathstudent55):

Let's use the first one as an example. Use a calculator for the divisions. 1/2 = 0.5 2/4 = 0.5 4/8 = 0.5 All divisions equal 0.5, and are equal, so the first one is a geometric series.

OpenStudy (mathstudent55):

Now do the same for the others.